# Meisam Kabiri1, and Holger Voos12 1 Interdisciplinary Center for Security Reliability and Trust (SnT)

### v3

@@ Abstract
This paper presents a novel approach that integrates 5G Time of Arrival (ToA) measurements into ORB-SLAM3 to enable global localization and enhance mapping capabilities for indoor drone navigation. We extend ORB-SLAM3’s optimization pipeline to jointly process ToA data from 5G base stations alongside visual and inertial measurements, while estimating system biases. This integration transforms the inherently local SLAM estimates into globally referenced trajectories and effectively resolves scale ambiguity in monocular configurations. Our method is evaluated using five real-world indoor datasets collected with RGB-D cameras and inertial measurement units (IMUs), complemented by simulated 5G ToA measurements at 28 GHz and 78 GHz frequencies using MATLAB and QuaDRiGa. Extensive experiments across four SLAM configurations (RGB-D, RGB-D-Inertial, Monocular, and Monocular-Inertial) demonstrate that ToA integration enables consistent global positioning across all modes while significantly improving local accuracy in minimal sensor setups. Notably, ToA-enhanced monocular SLAM achieves superior local accuracy (6.3 cm average) compared to the RGB-D baseline (11.5 cm), and enables reliable operation of monocular-inertial SLAM in scenarios where the baseline system fails completely. While ToA integration offers limited local accuracy improvements for sensor-rich configurations like RGB-D SLAM, it consistently enables robust global localization.
@@ I Introduction
The advent of 5G technology has introduced advanced positioning capabilities, particularly through Time of Arrival (ToA) measurements [10, 11]. ToA provides highly accurate distance estimates between receivers and known base station positions, offering centimeter-level precision in ideal conditions. By incorporating ToA data, SLAM systems can align their local frames with a global reference frame defined by the fixed positions of 5G base stations. This global alignment enables robust localization and mapping in GPS-denied environments, enhancing applications like inventory management, real-time monitoring, autonomous vehicles, augmented reality, and emergency response. Drones, for instance, can autonomously navigate warehouses, accurately mapping the environment and tracking inventory levels, as illustrated in Figure 1. By leveraging ToA measurements from two 5G base stations as global anchors, drones achieve higher accuracy and robustness, optimizing inventory management and real-time monitoring processes.
Figure 1: Illustration of an indoor warehouse environment where a drone navigates while two 5G base stations provide Time-of-Arrival (ToA) measurements. The base stations assist in refining the drone’s positional accuracy, supporting real-time localization and mapping (The warehouse figure was generated using OpenAI’s ChatGPT).
This paper extends ORB-SLAM3 [12] by integrating 5G ToA measurements into its optimization pipeline to achieve globally consistent SLAM. Specifically, we introduce an SE3 transformation node into the estimation process, deeply integrating it into both the front-end and back-end modules of ORB-SLAM3. This SE3 node estimates the rigid body transformation between the local SLAM reference frame and the global 5G-based reference frame in real-time, aligning the local map with the global coordinate system to ensure precise localization and mapping across different environments.
@@
We assess the effectiveness of our approach using five indoor datasets captured with RGB-D cameras and Inertial Measurement Units (IMUs). These datasets are further enhanced with simulated 5G Time of Arrival (ToA) measurements at frequencies of 28 GHz and 78 GHz, generated using MATLAB and QuaDRiGa. The experimental results indicate that our method maintains high local estimation precision, benefiting from the accuracy inherent in RGB-D-based ORB-SLAM3. Moreover, the incorporation of ToA measurements significantly strengthens global localization and mapping reliability. Notably, in scenarios where the conventional mono-inertial ORB-SLAM3 fails to sustain tracking, our method consistently delivers accurate localization throughout the entire trajectory.
@@ I-A Contributions
  • Integration of 5G ToA Measurements for Global SLAM: We introduce a novel approach that integrates 5G Time of Arrival (ToA) measurements into the ORB-SLAM3 framework. This integration anchors the SLAM system’s local reference frame to a global coordinate system defined by 5G base stations, enabling accurate and consistent global localization and mapping.
@@
  • Enhanced Robustness in Challenging Scenarios: The deep integration of ToA measurements into ORB-SLAM3’s optimization pipeline enhances the system’s robustness, particularly in environments where standard SLAM may struggle, such as feature-sparse or dynamic settings. Our approach maintains accurate localization even when standard mono-inertial ORB-SLAM3 loses tracking.
  • Comprehensive Evaluation Using Real and Simulated Datasets: We validate our approach using five real-world indoor datasets collected with RGB-D cameras and Inertial Measurement Units (IMUs), augmented with 5G ToA measurements simulated using MATLAB and QuaDRiGa. The evaluation demonstrates the effectiveness of our method in improving localization accuracy and robustness across various scenarios.
This paper is organized as follows. Section II reviews related work in visual SLAM, visual-inertial SLAM, and radio-based localization and SLAM. Section III details the methodology for integrating ToA measurements into ORB-SLAM3. Section IV describes the experimental setup and datasets used for evaluation. Section V presents and analyzes the results of the proposed method. Section VI discusses limitations and potential areas for future research, and Section VII concludes the paper with key findings and implications.
@@ II-C Radio-Based Localization and SLAM
To the best of our knowledge, this is the first study to combine 5G ToA measurements with visual-inertial SLAM for indoor drone navigation. Our approach addresses the limitations of prior methods by:
@@
  • Practical Implementation and Evaluation: Our work includes the practical integration of 5G Time of Arrival (ToA) measurements into the SLAM system and extensive evaluation using real-world indoor datasets augmented with simulated ToA data. We generate 5G ToA measurements using QuaDRiGa and validate our approach across multiple real-world datasets, demonstrating its effectiveness in realistic scenarios.
@@ III Methodology
Figure 2 illustrates the distinction between the local and global frames and highlights the transformation process that bridges them. The local map, represented in red, depicts the environment relative to the initial sensor position. However, this map lacks global context, making it unsuitable for applications requiring absolute positioning. Conversely, the global map, depicted in teal, integrates ToA measurements from fixed base stations, which serve as global reference points. The transformation $Tr$, shown in the figure, aligns the local frame to the global coordinate system, linking the two maps.
This alignment process is essential for achieving a globally consistent SLAM system. The transformation $Tr$ is estimated by leveraging ToA measurements, which provide spatial constraints based on the distances between the robot and the base stations. These constraints refine the local-to-global alignment and ensure that the trajectory and map points are accurately represented within the global frame. The dashed arrow in Figure 2 illustrates the influence of ToA measurements in bridging the local and global scales.
@@ III-B ToA Factor Formulation
ToA factor integrates distance measurements from base stations into the SLAM framework through a formulated error model. The factor connects multiple optimization vertices: the camera pose, local-to-global transformation, ToA measurement bias, and, in monocular cases, a scale factor. Figure 3 illustrates the structure of the ToA factor, highlighting how each component interacts within the optimization graph. The double-bordered base station nodes indicate that these parameters are fixed and their positions are known.
@@ III-C1 System Overview
Figure 4 illustrates the system architecture, highlighting how ToA measurements are incorporated into the SLAM pipeline. Several existing components have been modified, including the Local Bundle Adjustment module, which now incorporates ToA measurements for improved optimization accuracy. The system diagram uses color coding to distinguish between modifications to the original ORB-SLAM3 framework: new components are highlighted in green, while modified components are shown in yellow.
@@ III-C2 System Integration
The ToA measurements are integrated at multiple levels within the ORB-SLAM3 framework to maximize their utility while maintaining real-time performance, i.e., Tracking, Local Mapping, Loop Closing & ToA-based global map refinement. In what follows, we highlight some of the key changes and additions.
@@
Figure 5(d) illustrates the optimization graph structure. It comprises keyframe pose nodes, transformation nodes, and bias nodes, interconnected by three types of edge types, Odometry edges (straight lines), Covisibility edges (angled lines connecting from above or below), and ToA-related edges.
@@
The optimization graph for this transformation refinement process maintains a straightforward structure, incorporating only ToA and odometer edges. Figure 5(e) illustrates this graph structure.
Scale Refinement for Monocular Systems: Monocular visual SLAM systems inherently suffer from scale ambiguity. To address this, we incorporated a dedicated optimization component for global scale factor estimation and update.
This scale refinement is performed in addition to the local-to-global transformation while keeping keyframe poses fixed. By simplifying the graph to include only ToA edges, we enable global optimization, leveraging information from the entire keyframe map for a more accurate scale estimate. Figure 5(f) illustrates the structure of the scale refinement optimization graph. Key elements include: Keyframe pose nodes, scale factor node (s), ToA biases, and the Transformation node.
@@ IV Experiment
This section presents a detailed evaluation of the proposed methodology. Our experiments are designed to assess the performance of the SLAM system under diverse configurations, focusing on integrating simulated 5G ToA measurements with traditional visual and inertial data. The results provide insights into the impact of ToA integration across multiple datasets, sensor configurations, and 5G frequency bands.
@@ IV-A Experimental Setup and Data Collection
The experiments were conducted in a controlled indoor environment, the Aerolab, specifically designed to support drone flights and data collection under realistic conditions for SLAM evaluation. The core equipment included an Intel RealSense D435i camera, used for RGB-D (color and depth) and IMU data acquisition, and the OptiTrack motion capture system, which provided high-precision ground truth data through a 12-camera configuration. Integrating these components enabled a comprehensive setup for acquiring synchronized data essential for SLAM.
The Intel RealSense D435i camera, mounted on the drone, captured RGB-D data at 30 Hz and IMU data at 200 Hz, with automatic synchronization between the visual and inertial data streams. To further support SLAM evaluation, five unique datasets were collected by flying the drone through varied trajectories that mimic realistic navigation scenarios, capturing a wide range of motion. Additionally, simulated 5G ToA data was generated at a rate of 10 Hz to complement the visual and inertial data. OptiTrack ground truth data was collected at 120 Hz, ensuring high accuracy and alignment with the other sensor modalities.
The drone operated within a netted 5x5x5 meter flight area for safety, though the Aerolab itself is larger, providing additional space for equipment and experimental setups. Each dataset represents a different flight path and duration, as detailed below:
  • Dataset 0: 116 seconds
  • Dataset 1: 110 seconds
  • Dataset 2: 135 seconds
  • Dataset 3: 124 seconds
  • Dataset 4: 139 seconds
The trajectories for each dataset are visualized in Figure 6, where the top row displays the x-y trajectories and the bottom row illustrates the time vs. z plots.
(a) Dataset0
(b) Dataset1
(c) Dataset2
(d) Dataset3
(e) Dataset4
@@
The setup, including the drone equipped with the RealSense camera and the Aerolab environment, is illustrated in Figure 7. This figure provides visual context for the controlled flight space and the comprehensive equipment layout used for data collection.
The experiments were performed on a Ubuntu 20.04 laptop equipped with an Intel(R) Core(TM) i9-10885H CPU @ 2.40GHz with 16 cores and 32 GB of RAM.
To evaluate system robustness, we introduced significant noise to initial transformation estimates: 0.5m translation and $30^{\circ}$ rotation for MONO mode, and 1m translation with $90^{\circ}$ rotation for other modes. These substantial uncertainties effectively test the system’s trajectory recovery capabilities, simulating real-world scenarios where initial estimates may be imprecise.
@@
Figure 7: (a) Drone equipped with an Intel RealSense D435i camera for RGB-D and IMU data acquisition. (b) Experimental flight area within the netted environment for safe indoor drone testing. (c) Overall view of the Aerolab, illustrating both the drone testing area and additional laboratory equipment and space.
IV-B Calibration and Data Synchronization
The intrinsic and extrinsic calibration of the camera and IMU was conducted using the Kalibr toolbox and allan_variance_ros. The calibration procedure involved collecting data with checkerboard patterns and performing specific drone motions. This process resulted in intrinsic parameters such as focal length and distortion coefficients, as well as the transformation matrix between the camera and IMU, crucial for accurate sensor fusion. Data synchronization is essential for optimal SLAM performance. Since the RealSense camera provides both RGB-D and IMU data, these streams are automatically synchronized. The OptiTrack system’s ground truth data is also time-aligned with the collected data. The 5G ToA measurements are simulated based on the OptiTrack trajectories, ensuring that all datasets are synchronized.
IV-C Augmenting the Dataset with Simulated 5G ToA Measurements
Figure 8: Architecture of the dataset generation pipeline, showing the integration of simulated 5G ToA measurements with the Aerolab dataset. The pipeline consists of a 5G signal simulator using QuaDRiGa channel models, signal processing components (correlator and peak detection), and the final integration with IMU and image data from the Aerolab dataset. The combined data serves as input to the SLAM framework, with ToA measurements scaled by the speed of light to obtain distance measurements.
Acquiring real 5G ToA data in our lab setting is constrained by infrastructure and hardware limitations. To overcome this barrier, we employ a simulation-based approach to generate realistic ToA measurements. This strategy enables us to assess the advantages of integrating 5G positioning signals into the SLAM framework without the need for extensive physical network setups. By combining simulated ToA data with authentic drone flight data that includes RGB-D images, IMU readings, and precise ground truth trajectories, we construct an augmented dataset suitable for SLAM performance assessment. Figure 8 illustrates the overall structure of the simulation environment and how its components interact to generate this augmented dataset.
IV-D 5G ToA Simulation Methodology
Following our previous work [39, 38], we employ the same rigorous methodology to simulate 5G ToA measurements for the current dataset under the line-of-sight condition. The simulation framework utilizes the MATLAB 5G Toolbox to generate 5G signals, including Positioning Reference Signals (PRS) and Physical Downlink Shared Channel (PDSCH) resources, with consistent transmission parameters (0 dBm transmit power and 10 dB SNR).
The wireless channel characteristics are simulated using QuaDRiGa (quasi-deterministic radio channel generator) [40] with two distinct 5G FR2 (millimeter wave) network configurations: a 28 GHz configuration utilizing 200 MHz channel bandwidth with 120 kHz subcarrier spacing, and a 78 GHz configuration employing 400 MHz channel bandwidth with 240 kHz subcarrier spacing, both aligning with 3GPP specifications for FR2 deployments.
While the simulation framework remains unchanged, we adapt the base station placement to suit the current dataset’s environment. We virtually position four BSs within the experimental space, with coordinates defined in the OptiTrack system’s global frame as: BS1: $(2.5,-2.5,4.5)$, BS2: $(2.5,2.5,4.0)$, BS3: $(-2.5,2.5,5.0)$, BS4: $(-6.5,-2.5,2.0)$. The receiver is also assumed to be mounted on top of the drone.
The simulation maintains the established configuration parameters:
  • Signal generation at 10 Hz (every 0.1 seconds)
  • Empirically validated correlation threshold of 0.2 for ToA estimation
  • Omnidirectional antenna configurations for both transmitters and receivers
  • Consistent 5G system parameters across scenarios (bandwidth, subcarrier spacing, resource blocks, comb size, and cyclic prefix type)
The trajectory data from the current dataset is processed by computing velocities from consecutive pose measurements to maintain compatibility with the QuaDRiGa channel simulator’s requirements. This ensures that the simulated ToA measurements accurately reflect the dynamic characteristics of the platform’s motion in the new environment while maintaining consistency with our validated simulation methodology. The final statistics of the simulated ToA are given in Table I.

### v5

@@ Abstract
This paper presents a novel approach that integrates 5G Time of Arrival (ToA) measurements into ORB-SLAM3 to enable global localization and enhance mapping capabilities for indoor drone navigation. We extend ORB-SLAM3’s optimization pipeline to jointly process ToA data from 5G base stations alongside visual and inertial measurements while estimating system biases. This integration transforms the inherently local SLAM estimates into globally referenced trajectories and effectively resolves scale ambiguity in monocular configurations. Our method is evaluated using both Aerolab indoor datasets with RGB-D cameras and the EuRoC MAV benchmark, complemented by simulated 5G ToA measurements at 28 GHz and 78 GHz frequencies using MATLAB and QuaDRiGa. Extensive experiments across multiple SLAM configurations demonstrate that ToA integration enables consistent global positioning across all modes while maintaining local accuracy. For monocular configurations, ToA integration successfully resolves scale ambiguity and improves consistency. We further investigate scenarios with unknown base station positions and demonstrate that ToA measurements can effectively serve as an alternative to loop closure for drift correction. We also analyze how different geometric arrangements of base stations impact SLAM performance. Comparative analysis with state-of-the-art methods, including UWB-VO, confirms our approach’s robustness even with lower measurement frequencies and sequential base station operation. The results validate that 5G ToA integration provides substantial benefits for global SLAM applications, particularly in challenging indoor environments where accurate positioning is critical.
@@ I Introduction
The advent of 5G technology has introduced advanced positioning capabilities, particularly through ToA measurements [10, 11]. ToA provides highly accurate distance estimates between receivers and known base station positions, offering centimeter-level precision in ideal conditions. By incorporating ToA data, SLAM systems can align their local frames with a global reference frame defined by the fixed positions of 5G base stations. This global alignment enables robust localization and mapping in GPS-denied environments, enhancing applications like inventory management, real-time monitoring, autonomous vehicles, augmented reality, and emergency response. Drones, for instance, can autonomously navigate warehouses, accurately mapping the environment and tracking inventory levels, as illustrated in Fig. 1. By leveraging ToA measurements from 5G base stations as global anchors, drones achieve higher accuracy and robustness, optimizing inventory management and real-time monitoring processes.
Figure 1: Illustration of an indoor warehouse environment where a drone navigates while two 5G base stations provide ToA measurements. The base stations assist in refining the drone’s positional accuracy, supporting real-time localization and mapping (The warehouse figure was generated using OpenAI’s ChatGPT).
This paper extends ORB-SLAM3 [12] by integrating 5G ToA measurements into its optimization pipeline to achieve globally consistent SLAM. Specifically, we introduce an SE3 transformation node into the estimation process, deeply integrating it into both the front-end and back-end modules of ORB-SLAM3. This SE3 node estimates the rigid body transformation between the local SLAM reference frame and the global 5G-based reference frame in real-time, aligning the local map with the global coordinate system to ensure precise localization and mapping across different environments. Furthermore, we investigate challenging scenarios with unknown base station positions and limited visibility, demonstrating the system’s adaptability to real-world deployment constraints through sequential base station operation.
@@
We assess the effectiveness of our approach using five indoor datasets captured with RGB-D cameras and Inertial Measurement Units (IMUs) from our Aerolab facility, as well as the standard EuRoC MAV benchmark for comparison with state-of-the-art methods. These datasets are further enhanced with simulated 5G ToA measurements at frequencies of 28 GHz and 78 GHz, generated using MATLAB and QuaDRiGa. The experimental results indicate that our method maintains high local estimation precision, benefiting from the accuracy inherent in ORB-SLAM3. Moreover, the incorporation of ToA measurements enables global localization and mapping. We also demonstrate that ToA integration can effectively serve as an alternative to loop closure mechanisms for drift correction, maintaining trajectory consistency even in challenging scenarios where loop detection fails or is unavailable. Comparative analysis with methods like UWB-VO confirms the robustness of our approach even with lower measurement frequencies and more realistic operational constraints. Finally, we evaluate and quantify how different geometric arrangements of base stations (Tetrahedral, Z-Shape, Asymmetric, Diamond, and Clustered) affect SLAM performance, demonstrating that optimal configurations can improve localization accuracy.
@@ I-A Contributions
  • Integration of 5G ToA Measurements for Global SLAM: We introduce a novel approach that integrates 5G ToA measurements into the ORB-SLAM3 framework. This integration anchors the SLAM system’s local reference frame to a global coordinate system defined by 5G base stations, enabling accurate and consistent global localization and mapping.
@@
  • Enhanced Robustness in Challenging Scenarios: We demonstrate system performance with unknown base station positions and sequential visibility constraints, showing that even partial or intermittent ToA measurements can significantly improve localization accuracy and consistency.
  • ToA as an Alternative to Loop Closure: We evaluate the effectiveness of ToA integration as a substitute for traditional loop closure mechanisms, showing comparable or superior performance in maintaining trajectory consistency, particularly beneficial in environments where loop detection is unreliable or impossible.
  • Comprehensive Evaluation Against State-of-the-Art: We validate our approach using multiple datasets including both custom Aerolab collections and the standard EuRoC MAV benchmark, providing direct comparisons with state-of-the-art methods like UWB-VO under realistic operational constraints.
This paper is organized as follows. Section II reviews related work in visual SLAM, visual-inertial SLAM, and radio-based localization and SLAM. Section III details the methodology for integrating ToA measurements into ORB-SLAM3. Section IV describes the experimental setup and datasets used for evaluation. Section V presents and analyzes the results of the proposed method. Section VII discusses limitations and potential areas for future research, and Section VIII concludes the paper with key findings and implications.
@@ II-C Radio-Based Localization and SLAM
The integration of Ultra-Wideband (UWB) technology with visual SLAM has emerged as another promising approach for improving localization accuracy. Lin and Yeh [40] proposed a drift-free visual SLAM technique for mobile robot localization by integrating UWB positioning technology. Their approach utilizes the global constraint of UWB positioning to reduce locally accumulated errors of visual SLAM localization based on the extended Kalman filtering (EKF) framework. Experimental results demonstrated that the integration of UWB positioning reduced the overall drift error of robot navigation by more than 50%. However, this approach relies on a dense network of UWB anchors and primarily addresses drift reduction rather than providing a comprehensive framework for global localization.
Li et al. [41] proposed UWB-VO, a tightly coupled SLAM framework that combines monocular vision with UWB distance measurements. Their method jointly optimizes UWB distance residuals and visual reprojection errors to improve positioning accuracy and address scale ambiguity in monocular SLAM, achieving scale error rates below 1%. The system features a UWB anchor position optimization technique for rapid initialization and provides comparative benchmarks against several visual-inertial odometry algorithms. However, UWB-VO operates under the assumption of continuous visibility of at least one UWB anchor and requires a relatively high measurement frequency (30 Hz), which may limit its practicality in dynamic or complex environments.
Research on alternatives to traditional loop closure for drift correction represents another important direction in SLAM development. Conventional loop closure mechanisms rely on revisiting previously mapped areas, which may not be feasible in exploration scenarios, time-critical operations, or environments with significant appearance changes. While several approaches have explored external reference sources for drift correction, few have systematically evaluated radio-based measurements as direct substitutes for loop closure. This gap highlights the need for robust SLAM systems that can maintain trajectory consistency even with limited or intermittent external measurements from imprecisely localized reference points.
@@
To the best of our knowledge, this is the first study to comprehensively evaluate 5G ToA integration with visual-inertial SLAM across both known and unknown base station scenarios. Our approach addresses the limitations of prior methods by:
@@
  • Robustness to Unknown Base Station Configurations: We demonstrate system performance with unknown base station positions and sequential visibility, showing that our approach remains effective even with partial or intermittent ToA measurements—conditions commonly encountered in real-world deployments.
  • Alternative to Traditional Loop Closure: We provide the first systematic evaluation of ToA integration as a substitute for conventional loop closure mechanisms, demonstrating comparable or superior drift correction without requiring revisits to previously mapped areas.
  • Practical Implementation with Realistic Constraints: Unlike approaches such as UWB-VO that require high-frequency measurements (30 Hz) and continuous anchor visibility, our system operates effectively with lower measurement rates (5-10 Hz) and intermittent base station connectivity, making it more practical for real-world deployment.
@@ III Methodology
Fig. 2 illustrates the distinction between the local and global frames and highlights the transformation process that bridges them. The local map, represented in red, depicts the environment relative to the initial sensor position. However, this map lacks global context, making it unsuitable for applications requiring absolute positioning. Conversely, the global map, depicted in teal, integrates ToA measurements from fixed base stations, which serve as global reference points. The transformation $Tr$, shown in the figure, aligns the local frame to the global coordinate system, linking the two maps.
This alignment process is essential for achieving a globally consistent SLAM system. The transformation $Tr$ is estimated by leveraging ToA measurements, which provide spatial constraints based on the distances between the robot and the base stations. These constraints refine the local-to-global alignment and ensure that the trajectory and map points are accurately represented within the global frame. The dashed arrow in Fig. 2 illustrates the influence of ToA measurements in bridging the local and global scales.
@@ III-B ToA Factor Formulation
ToA factor integrates distance measurements from base stations into the SLAM framework through a formulated error model. The factor connects multiple optimization vertices: the camera pose, local-to-global transformation, ToA measurement bias, and, in monocular cases, a scale factor. Fig. 3 illustrates the structure of the ToA factor, highlighting how each component interacts within the optimization graph. The double-bordered base station nodes indicate that these parameters are fixed and their positions are known.
@@ III-C1 System Overview
Fig. 4 illustrates the system architecture, highlighting how ToA measurements are incorporated into the SLAM pipeline. Several existing components have been modified, including the Local Bundle Adjustment module, which now incorporates ToA measurements for improved optimization accuracy. The system diagram uses color coding to distinguish between modifications to the original ORB-SLAM3 framework: new components are highlighted in green, while modified components are shown in yellow.
@@ III-C2 System Integration
The ToA measurements are integrated at multiple levels within the ORB-SLAM3 framework to maximize their utility while maintaining real-time performance, i.e., Tracking, Local Mapping, Loop Closing-ToA-based global map refinement. In what follows, we highlight some of the key changes and additions.
@@
Fig. 5(d) illustrates the optimization graph structure. It comprises keyframe pose nodes, transformation nodes, and bias nodes, interconnected by three types of edge types, Odometry edges (straight lines), Covisibility edges (angled lines connecting from above or below), and ToA-related edges.
@@
The optimization graph for this transformation refinement process maintains a straightforward structure, incorporating only ToA and odometer edges. Fig. 5(e) illustrates this graph structure.
Scale Refinement for Monocular Systems: Monocular visual SLAM systems inherently suffer from scale ambiguity. To address this, we incorporated a dedicated optimization component for global scale factor estimation and update (Fig. 5(f)).
This scale refinement is performed in addition to the local-to-global transformation while keeping keyframe poses fixed. By simplifying the graph to include only ToA edges, we enable global optimization, leveraging information from the entire keyframe map for a more accurate scale estimate. Fig. 5(f) illustrates the structure of the scale refinement optimization graph. Key elements include: Keyframe pose nodes, scale factor node (s), ToA biases, and the Transformation node.
@@
III-D System Operation with Unknown Base Station Positions
While our primary implementation assumes known base station locations to enable global positioning, we also investigate scenarios where these positions are unknown. In this configuration, we modify our optimization framework by changing the base station nodes from fixed to non-fixed status while maintaining a fixed transformation node (TR). This adjustment effectively eliminates the requirement for prior knowledge of base station locations.
The modified optimization graph structure still incorporates ToA measurements, but instead of using them to align the local frame with a globally defined coordinate system, they now serve primarily to constrain the relative geometry of the trajectory and enhance local consistency. While this configuration precludes absolute global positioning capabilities, it remains valuable for improving SLAM performance, particularly in minimal sensor setups where scale ambiguity and drift represent significant challenges.
@@ IV Experiment
This section presents a detailed evaluation of the proposed methodology, assessing SLAM performance under diverse configurations with simulated 5G ToA measurements.
@@ IV-A Experimental Setup and Data Collection
The experiments were conducted using two complementary datasets:
Aerolab Datasets: Our custom datasets were collected in the Aerolab indoor environment, designed for drone flights under realistic conditions. We used an Intel RealSense D435i camera for RGB-D (30 Hz) and IMU (200 Hz) data acquisition, with an OptiTrack motion capture system (12 cameras at 120 Hz) providing high-precision ground truth. Five unique datasets were collected through varied flight trajectories ranging from 110-139 seconds Fig. 6. The drone operated within a netted 5x5x5 meter flight area, as shown in Fig. 7. Calibration of the RealSense camera and IMU was performed using the Kalibr toolbox and allan_variance_ros, with data streams automatically synchronized. The simulated ToA measurements were time-aligned with the collected data based on the OptiTrack trajectories.
EuRoC MAV Dataset: To enable comparison with state-of-the-art methods, we also evaluated our approach on the public EuRoC MAV benchmark [42]. This dataset provides stereo images, IMU measurements, and Vicon motion capture ground truth across various indoor environments with different complexity levels. For the Machine Hall sequences (MH01-MH05), we used simple Gaussian ToA simulation to match the conditions in UWB-VO experiments. For the Vicon Room sequences (V101-V203), we employed our comprehensive 5G ToA simulation methodology as detailed in our previous work [38, 39].
The experiments were performed on a Ubuntu 20.04 laptop with an Intel i9-10885H CPU and 32 GB of RAM.
(a) AL01
(b) AL02
(c) AL03
(d) AL04
(e) AL05
@@
Figure 7: (a) Drone equipped with Intel RealSense D435i camera. (b) Experimental flight area within netted environment. (c) Overall Aerolab view.
IV-B 5G ToA Simulation and Data Integration
Due to infrastructure and hardware constraints, we employed a simulation-based approach to generate realistic ToA measurements. The simulation framework, illustrated in Fig. 8, utilizes MATLAB 5G Toolbox for signal generation and QuaDRiGa [43] for channel modeling, with two distinct 5G FR2 configurations:
  • 28 GHz: 200 MHz bandwidth, 120 kHz subcarrier spacing
  • 78 GHz: 400 MHz bandwidth, 240 kHz subcarrier spacing
Figure 8: Architecture of the dataset generation pipeline, showing the integration of simulated 5G ToA measurements with the Aerolab dataset. The combined data serves as input to the SLAM framework.
Four virtual base stations were positioned within the experimental space at coordinates BS1: $(2.5,-2.5,4.5)$, BS2: $(2.5,2.5,4.0)$, BS3: $(-2.5,2.5,5.0)$, and BS4: $(-6.5,-2.5,2.0)$ in the OptiTrack’s global frame. Signals were generated at 10 Hz with consistent parameters (0 dBm transmit power, 10 dB SNR) and omnidirectional antennas.